OpenStudy (anonymous):

61 62 63 (4+ 4 +4+ ) is divisible by

OpenStudy (shayaan_mustafa):

kindly make this question more clear.

OpenStudy (anonymous):

that means 4 squre root of 61 62 like that

OpenStudy (shayaan_mustafa):

what are plus sign in mid ? why is there gap??

OpenStudy (anonymous):

\[the greatest among the numbers \sqrt[4]{3}\sqrt[5]{4}\sqrt[10]{12}1\]